# Surf-NeRF: Surface Regularised Neural Radiance Fields

### v1

@@ 3.4 Local Smoothness
The spatial sampling batch encompasses the region in front of and behind the candidate surface. Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance $\quantity|\frac{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}{\norm{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}}\cdot\hat{\mathbf{n}}_{\star}|$ from the surface encouraging a plane of density to form perpendicular to the normal. Points far from the surface should have low density, whilst those on the same plane as the sample point should be dense. This loss term,
@@ 3.5 Visually Realistic Appearance Separation
Whilst highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity. This is characterised by incorrect geometry with a view-dependent colour bias; that is the minimum specular colour over all viewing angles is non-zero to satisfy the appearance in all training view-points. More details are found in the supplementary.
@@
is based upon the $k$-Nearest Neighbours of each sample point and weighted by the cosine distance between the sample and its neighbours. This approach provides a localised consistency which is edge-aware continuous over the sampling sphere, capturing occlusions. Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model and are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
@@ 3.6 Curriculum Learning
We impose our surface regularisation under curriculum learning to maintain visual fidelity whilst still improving geometry. As training progresses and additional regularisation is imposed more frequently, the representation is refined to satisfy the geometric and appearance characteristics of a surface light field. The surface constraints presented in Sections 3.4-3.5 are of similar computational complexity to a regular training step, requiring a second pass through both MLPs and therefore taking close to double the time. By using curriculum learning we trade-off the improvement to the representation with training time.
@@ 4 Results
Our regularisation is implemented in JAX [13] and based on the ZipNeRF codebase [7]. We utilise the improvements to quality and speed in ZipNeRF combined with the physics-based model structure introduced in Ref-NeRF. We compare to prior works MipNeRF360 [4] and Ref-NeRF [36]. We evaluate Surf-NeRF on the Shiny Real and the Shiny Objects [36] datasets used in Ref-NeRF as the main benchmark for our approach. We also evaluate on a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset. Each scene contains approximately 40 training images and 10 test images with ground truth poses obtained using a Universal Robots UR5e robotic arm. We select objects that include non-planar specularities, fine details and specularities which present as virtual images in front of the surface. Further details are provided in the supplementary material.
@@ 4.1 Shiny Objects Dataset
Table 1 summarises results on the Shiny Objects dataset, comparing both visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of the density derived normals in degrees and root mean squared error (RMSE) of the disparity in inverse units. We compare across positionally encoded baselines based on MipNeRF360 [4], denoted as “PE-based” and “grid-based” implementations based on ZipNeRF [6]. We also compare against a variant of MipNeRF which includes a diffuse colour channel, denoted as “MipNeRF+Diff”. We train the positionally encoded implementations using V2-8 TPUs, and the grid-based methods on four NVIDIA V100 32GB GPUs. Per scene results are included in the supplementary materials.
Table 1: Summary on the Shiny Objects [36] dataset between visual metrics (PSNR, SSIM) and geometric metrics (MAE, RMSE). Our method yields more accurate normals, trading off a maginal decrease to PSNR. Yellow, orange and red are the third, second, and first scores.
@@
[…]
Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity whilst encouraging a more geometrically accurate representation. In the positionally encoded approaches we see a 14.4% improvement to normal accuracy, with notable improvements around regions of high specularity. We also see a 9.2% improvement to normals for the grid based implementation. Given the discretisation of the grid-based approaches, we see comparatively higher disparity errors across the dataset but improved normals from the interpolation of encodings. Our approach realises a 6.7% improvement in disparity versus Zip+Ref-NeRF.
In Figure 4 we demonstrate that the proposed regularisation is able to correct for warping of geometry and improve its consistency on the car scene, whilst also correctly separating the Lambertian racing stripes. Similarly, on the toaster dataset we successfully separate the Lambertian toast whilst retaining its reflection and improving geometry around specular materials with high curvature.
Figure 4: Results from the grid based models. By regularising the scene, we not only remove Lambertian scene content from the specular, like the stripes on the car and toast in the toaster, but our approach yields improved geometry particularly around curves and specularity, like the car windshield and corners of the toaster. Ref-NeRF* indicates the Zip+Ref-NeRF model.
Figure 5 depicts performance of positionally encoded models on specular surfaces with low texture. Surf-NeRF yields drastically better surface geometry and normals.
Figure 5: Geometry results on the non-grid based models. The surface regularisation can resolve errors in geometry which stem from shape-radiance ambiguity e.g. in regions with low-frequency textures on the teapot lid, without compromising visual fidelity. This attains substantially improved normals in these regions.
@@ 4.2 Shiny Real Dataset
The visual fidelity results on the Shiny Real dataset [36] are shown in Table 2 for Surf-NeRF and comparison baseline models. We demonstrate comparable visual fidelity to other state of the art NeRF variants demonstrating no appreciable drop in visual fidelity compared to non-regularised approaches on the positional encoding based models. The complex appearances of these datasets combined with the hash-encoding sees a small drop in PSNR values for grid-based variants. We show decreased reliance on floaters to explain appearance via Surf-NeRF in the supplement.
@@
[…]
Surf-NeRF (Ours) 23.26 0.573
[…]
@@ 4.3 Koala Dataset
Table 3 shows the strength of our proposed regularisation on complex reflective scenes. Figure 6 demonstrates the ability for the proposed regularisation to separate Lambertian and specular colours in a more physically consistent manner, and to overcome failure cases where specularities are not densely sampled by the training data. Additional results are provided in the supplementary.
@@
Surf-NeRF (Ours) 27.24 0.655
Figure 6: The “gold head” and “shiny ball” scenes from our captured Koala dataset. The first row contains RGB and median depth renderings, whilst the second contains the diffuse and specular renderings. The Gold Head dataset shows complete separation of Lambertian and specular appearance for the mannequin. The Shiny Ball dataset represents a failure case for the Zip+Ref-NeRF baseline, however the proposed regularisation regularises these floaters away during training and appropriately places the majority of the colour into the diffuse colour term.
@@ 4.4 Ablation Study
In Table 4 we present an ablation study of our additional surface regularisation terms on the Helmet scene from the Shiny Objects dataset. As shown, including our full approach provides strong performance across both the visual and geometric metrics. Our appearance based regularisation is coupled closely to the geometry of the scene. Removing the specular bias term drastically increases the normal error, however does not alter disparity in the scene. The normal term improves continuity of the geometry in the scene. Whilst removing it allows for each individual normal to be determined more freely, reducing the MAE, it also allows for floaters to be introduced unnecessarily by our spatial term corresponding to an increase in PSNR but also an increase in disparity error within the scene.
Table 4: Ablation study on Surf-NeRF regularisation terms on the Helmet scene. All terms contribute to overall performance.
@@
[…]
w/o spatial 32.33 0.972 10.67 0.196
[…]
Table 5: Parameter study on regularisation frequency on Coffee scene. There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1.
@@
1024 8 31.38 0.967 14.89 0.181
512 4 31.99 0.968 15.29 0.181
256 2 31.70 0.964 18.17 0.208
128 1 32.09 0.967 15.95 0.181
Table 5 demonstrates an study on our curriculum learning approach. We demonstrate the effectiveness of our approach as a function of the frequency of regularisation. We demonstrate improvement to the normals and disparity in all cases compared to Zip+Ref-NeRF. Note that performance is highest for regularisation at a rate which is sparse enough for adaptation to changing density and frequent enough that corrections to the representation are retained.
@@ 4.5 Surface Regularisation as Finetuning
We demonstrate surface regularisation as a fine-tuning step, after a NeRF has been trained on pre-trained Zip+Ref-NeRF and ZipNeRF model on the car scene from the Shiny Objects dataset. As ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, an example of our applicability to models without reflection parameterisation. The spatial sampling terms depend only on the geometry of the NeRF and so may be applied to other density-based neural fields without modification. We continue training at a fixed learning rate of $3.25\times 10^{-4}$ and present our results in Table 6. In both cases we demonstrate an improvement in normals over both the base model and the naïve approach where the model continues training without our regularisation. Note that finetuning with the reflection parameterisation, as in our Zip+RefNeRF model, results in higher disparity errors as the NeRF changes density placement during the additional training.
Table 6: Surf-NeRF as a finetuning step and applicability to other NeRF variants. With little additional training, we refine surface normals.
@@ 4.6 Limitations
Modelling the scene as surfaces, elements that are better represented volumetrically like hair or subsurface scattering may be unnecessarily regularised and their representation degraded using our proposed regularisation. Since Surf-NeRF makes use of a second pass through the network, our methodology takes twice as long on each regularisation step compared to a vanilla training step. We enforce geometric consistency during training at the cost of restricting the networks ability to deliver the best photometric accuracy, leading to marginally reduced PSNR scores.
@@ 5 Conclusion
We have proposed Surf-NeRF, a novel regularisation approach and four novel physically derived regularisers, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model. This ensures consistency of density and normals on geometry in the volume, in addition to a physics-inspired regularisation on view-dependent appearance. This enables a NeRF to adjust geometric structure during training without greatly sacrificing visual fidelity, demonstrating qualitatively improved continuity of geometry, with improvements of up to 14.4% in normals and 6.7% in disparity.
This work produces better NeRF representations for geometry critical tasks with unchanged model architectures.

### v2

@@ 3.4 Local Smoothness
The spatial sampling batch encompasses the region in front of and behind the candidate surface. Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance from the surface encouraging a plane of density to form perpendicular to the normal. Points far from the surface should have low density, and those on the same plane as the sample point should be dense. This loss term,
@@ 3.5 Visually Realistic Appearance Separation
Although highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity. This is characterised by incorrect geometry with a view-dependent colour bias; that is the minimum specular colour over all viewing angles is non-zero to satisfy the appearance in all training view-points. More details are found in the supplementary.
@@
is based upon the $k$-Nearest Neighbours of each sample point and weighted by the cosine distance between the sample and its neighbours. This approach provides a localised consistency which is edge-aware continuous over the sampling sphere, capturing occlusions. Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model. These are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
@@ 3.6 Curriculum Learning
We impose our surface regularisation under curriculum learning to maintain visual fidelity and improve geometry. As training progresses and additional regularisation is imposed more frequently, the representation is refined to satisfy the geometric and appearance characteristics of a surface light field. The surface constraints presented in Sections 3.4-3.5 are of similar computational complexity to a regular training step, requiring a second pass through both MLPs and therefore taking close to double the time. By using curriculum learning we trade off the improvement to the representation with training time.
@@ 4 Results
Our regularisation is implemented in JAX [13] and based on the ZipNeRF codebase [7]. We utilise the improvements to quality and speed in ZipNeRF combined with the physics-based model structure introduced in Ref-NeRF. We compare to prior works MipNeRF360 [5] and Ref-NeRF [43].
We evaluate Surf-NeRF on the Shiny Real and the Shiny Objects [43] datasets used in Ref-NeRF as the main benchmark for our approach. We also introduce a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset. Each scene contains approximately 40 training images and 10 test images with ground truth poses obtained using a Universal Robots UR5e robotic arm. We select objects that include non-planar specularities, fine details and specularities which present as virtual images in front of the surface. Further details are provided in the supplementary material.
@@ 4.1 Shiny Objects Dataset
Table 1 summarises results on the Shiny Objects dataset. We compare visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of surface normals in degrees and root mean squared error (RMSE) of the disparity in inverse units. We compare across positionally encoded methods (PE) based on MipNeRF360 [5] and “grid-based” methods (Grid) based on ZipNeRF [6]. This includes baselines: MipNeRF with a diffuse channel (MipNeRF+Diff), and lattice and cubic grid based ZipNeRF with reflection parameterisation (Zip+Ref-NeRF). We train the PE implementations using V2-8 TPUs, and the Grid methods on four NVIDIA V100 32GB GPUs. Scene results are included in the supplementary materials.
Table 1: Summary on the Shiny Objects [43] dataset between visual metrics (PSNR, SSIM) and geometric metrics (MAE, RMSE). Our method yields significantly more accurate normals, trading off a marginal decrease to PSNR. Yellow, orange and red are the third, second, and first scores.
@@
MipNeRF+Diff 30.84 0.936 60.00 0.122
Ref-NeRF 33.21 0.971 25.89 0.117
Surf-NeRF (Reg. Only) 33.01 0.967 22.15 0.117
Grid ZipNeRF 31.02 0.952 19.47 0.209
Zip+Ref-NeRF 33.14 0.964 14.70 0.195
Lattice+Ref-NeRF 33.21 0.964 13.76 0.202
Surf-NeRF (No Lattice) 32.13 0.954 13.35 0.202
Surf-NeRF (Ours) 32.82 0.955 10.60 0.198
Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity and encourage a more geometrically accurate representation. In the positionally encoded approaches we see a 14.4% improvement to normal accuracy, with notable improvements around regions of high specularity. Surf-NeRF attains a 27.9% improvement to normals for the grid based implementation. Given the discretisation of the grid-based approaches, we see comparatively higher disparity errors across the dataset but improved normals from the interpolation of encodings.
In Figure 4 we demonstrate that the proposed regularisation is able to substantially improve surface normals and the surface consistency. This also separates the Lambertian scene content, including the racing stripes as in 5. Similarly, on the toaster scene we successfully separate the Lambertian toast and retain its reflection in the specular term.
Figure 4: Geometry comparison between normals achieved by Surf-NeRF and traditional grid implementations on the car and coffee scenes. Ref-NeRF* indicates the Zip+Ref-NeRF model. Regularising the density field improves the continuity of surfaces reducing holes and divots. The inclusion of the permutohedral lattice encoding significantly improves surface normals by better approximating smooth surface curvature.
Figure 5: Appearance separation between Surf-NeRF and baseline method Zip+Ref-NeRF. Colour renderings are comparable, but Surf-NeRF produces physically based separation of Lambertian and specular scene content, such the racing stripes and toast.
@@ 4.2 Shiny Real Dataset
The visual fidelity results on the Shiny Real dataset [43] are shown in Table 2 for Surf-NeRF and comparison baseline models. We demonstrate comparable visual fidelity to other state of the art NeRF variants demonstrating no appreciable drop in visual fidelity compared to non-regularised approaches on the positional encoding based models. The complex appearances of these datasets combined with the hash encoding sees a small drop in PSNR values for grid-based variants. Surf-NeRF shows decreased reliance on floaters to explain appearance as shown in the supplement.
@@
Ref-NeRF 23.65 0.529
Surf-NeRF (Reg. Only) 23.26 0.573
Grid ZipNeRF 23.63 0.626
Zip+RefNeRF 22.63 0.626
Lattice+RefNeRF 22.94 0.614
Surf-NeRF (No Lattice) 22.48 0.600
Surf-NeRF (Ours) 23.04 0.600
@@ 4.3 Koala Dataset
Table 3 shows the strength of the proposed regularisation on complex reflective scenes. Figure 6 demonstrates the ability for the proposed regularisation to separate Lambertian and specular colours in a more physically consistent manner, and to overcome failure cases where specularities are not densely sampled by the training data. Additional results are provided in the supplementary.
@@
Surf-NeRF (No Lattice) 27.24 0.655
Surf-NeRF (Ours) 26.86 0.654
Figure 6: The “Gold Head” and “Shiny Ball” scenes from the Koala dataset. Top rows contain RGB and median depth renderings. The Gold Head scene shows complete separation of Lambertian and specular appearance for the mannequin. The Shiny Ball dataset represents a failure case for Zip+Ref-NeRF baseline, however the proposed regularisation successfully reconstructs and appropriately places the majority of the colour into the diffuse channel.
@@ 4.4 Ablation Study
As shown in Table 4 including our full approach provides strong performance across both the visual and geometric metrics on Shiny Objects [43]. Our appearance based regularisation is coupled closely to the geometry of the scene. The permutohedral lattice encoding significantly improves surface normals coupled with our regularisation.
Table 4: Ablation study on Surf-NeRF regularisation terms on the Shiny Objects dataset. All terms contribute to overall performance. $\bigstar$ ablates deterministic sphere sampling, # ablates the permutohedral lattice and all $\mathcal{L}$ terms match those in the paper.
@@
33.03 0.962 15.47 0.181
32.55 0.959 15.81 0.192
32.83 0.961 14.83 0.200
33.32 0.965 16.34 0.196
33.12 0.964 12.21 0.183
32.76 0.958 11.40 0.196
32.96 0.964 10.90 0.199
33.14 0.969 10.74 0.201
Rand. Samples 32.26 0.954 13.70 0.203
(No Lattice) 32.13 0.954 13.35 0.202
(Ours) 32.82 0.955 10.60 0.198
Table 5: Parameter study on regularisation frequency. There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1. The 512/4 regime balanced multi-view photometric losses with the surface light field regularisation.
@@
1024 8 32.63 0.955 12.46 0.204
512 4 32.82 0.953 10.60 0.198
256 2 30.59 0.946 18.99 0.220
128 1 29.77 0.926 28.37 0.218
Table 5 demonstrates a parameter study on our curriculum learning approach. Regularising frequency forces the NeRF to learn better separation of Lambertian and view-dependent channels, as in Figure 1 but reduces overall geometric quality by strongly manipulating density earlier in training. See the supplementary for more detail.
@@ 4.5 Surface Regularisation as Finetuning
We demonstrate surface regularisation as a fine-tuning step, after a NeRF has been trained on pre-trained Zip+Ref-NeRF and ZipNeRF model on the car scene from the Shiny Objects dataset. Since ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, demonstrating applicability to models without reflection parameterisation. The spatial sampling terms depend only on the geometry and can be applied to other density-based neural fields without modification. We continue training at a fixed learning rate of $3.25\times 10^{-4}$ and present our results in Table 6. Normals are improved over the base model and naïve approach where the model continues training without our regularisation. Finetuning with the reflection parameterisation, as in our Zip+RefNeRF model slightly increases disparity errors as the NeRF changes density placement during the additional training.
Table 6: Surf-NeRF as a finetuning step with applicability to other NeRF variants. Additional training refines surface normals.
@@ 4.6 Limitations
Modelling the scene as surfaces, volumetric elements like hair or subsurface scattering may have their representation degraded with the proposed regularisation. Our methodology takes twice as long on each regularisation step compared to a vanilla training step owing to two passes through the network. We enforce geometric consistency during training at the cost of reducing photometric accuracy, leading to slightly reduced PSNR scores.
@@ 5 Conclusion
We have proposed Surf-NeRF, a novel regularisation approach with four physically derived regularisation terms and a lattice hash encoding, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model. This ensures smoothly varying surface normals and geometry with more realistic Lambertian and view-dependent appearance. This enables a NeRF to adjust geometric structure without greatly sacrificing visual fidelity, improving accuracy of up to 27.8% in normals and 6.7% in disparity. This work is a key step towards better radiance-based representations for geometry critical tasks.